OpenStudy (asbhardwaj18):

If root(x) = t + 7 then which option is correct ? (i) v directly proportional to t (ii) v inversly proportional to t (iii) v directly proportional to root(t) (iv) a = constant

OpenStudy (mrnood):

there is no reference to v or a in the equation

OpenStudy (mrnood):

is 'x' intended to represent displacement? v to be velocity a to be acceleration?

OpenStudy (asbhardwaj18):

yes

OpenStudy (shamim):

U know velocity v=dx/dt

OpenStudy (shamim):

Acceleration a=dv/dt

OpenStudy (shamim):

Do u know hw to differentiate √x with respect to t

OpenStudy (shamim):

d√x/dt=?

OpenStudy (asbhardwaj18):

no

OpenStudy (shamim):

Ok.

OpenStudy (shamim):

U know the formula dx^3/dx=3*x^(3-1)=3x^2 Right?

OpenStudy (shamim):

Again dx^2/dx=2*x^(2-1)=2x Right?

OpenStudy (shamim):

d√x/dx=dx^(1/2)/dx=(1/2)*x^(-1/2)*dx/dt=?